Refer to the graphs. Assume that pizza is measured in slices and beer in pints. In which of the graphs is the opportunity cost of a pint of beer the lowest? There are four graphs. Graph A has points (0,8) and (5, 0). Graph B has points (0, 5) and (8,0), graph C has points (0, 1.25) and (2.00, 0), and graph d (0, 10) and (10,0). The "x" axis represents beer and the "y" axis represents pizza.
My work: I tried to add extra points on the graph to figure it out. Then I tried to compare the graphs and select a graph where there wasn't a wide distrance between the points on the x axis and y axis.

Answer by math_tutor2020(3835)


The graphs aren't shown. There's missing info.

For this problem I'll assume that each graph is linear.
If the assumption is incorrect, then please let me know so I can fix things.

x = amount of beer in pints
y = amount of pizza in number of slices
A point like (x,y) = (0,8) means x = 0 and y = 8 pair up together
We have 0 pints of beer and 8 slices of pizza.

Graph A goes through (0,8) and (5,0)
Compute the slope
m = slope
m = rise/run
m = (change in y)/(change in x)
m = (y2 - y1)/(x2 - x1)
m = (0-8)/(5-0)
m = -8/5
m = -1.6
The slope is m = -1.6

This tells us that each time x goes up by 1, then y goes down by 1.6
It means each time you buy a pint of beer, the number of slices you buy goes down by 1.6 slices.
In other words, for each pint of beer, you give up 1.6 slices of pizza.

We want to flip things so that we find out how much beer we give up for each extra slice purchased.
This means we'll need the reciprocal, which conveniently enough flips the fraction.

Apply the reciprocal
-8/5 flips to -5/8 = -0.625
Doing this let's us see that each time y goes up by 1, x will drop by 0.625
For each slice of pizza you get, you give up the chance of buying 0.625 pints of beer.
This is the opportunity cost for the beer in graph A.

Compute the slope of graph B
m = rise/run
m = (change in y)/(change in x)
m = (y2 - y1)/(x2 - x1)
m = (0-5)/(8-0)
m = -5/8
m = -0.625
Then apply the reciprocal
-5/8 flips to -8/5 = -1.6
For each slice of pizza you get, you give up 1.6 pints of beer. This result (1.6) is larger than the previous result (0.625)
Therefore graph A has the lowest opportunity cost so far.

Now find the slope of graph C
m = rise/run
m = (change in y)/(change in x)
m = (y2 - y1)/(x2 - x1)
m = (0-1.25)/(2-0)
m = -1.25/2
m = -0.625
m = -5/8
This flips to -8/5 = -1.6
We get the same opportunity cost as graph B.

Lastly compute the slope of graph D
m = slope
m = rise/run
m = (change in y)/(change in x)
m = (y2 - y1)/(x2 - x1)
m = (0-10)/(10-0)
m = -10/10
m = -1
The reciprocal of this is -1/1 or simply -1
Nothing changes.
For each slice of pizza you get, you give up 1 pint of beer.
This opportunity cost (1) is not smaller than graph A's opportunity cost (0.625)

Here is a summary of the opportunity costs of beer for each graph
  • A) 0.625
  • B) 1.6
  • C) 1.6
  • D) 1


Answer: Graph A has the lowest opportunity cost for the pints of beer.